OpenStudy (anonymous):

derivative of v(t)= t^2e^(-ct)

OpenStudy (anonymous):

use the product rule v' = t^2 * -c * e^(-ct) + e^(-ct)* 2t = te^(-ct)(2 - ct)

OpenStudy (anonymous):

how would you solve using the chain rule?

OpenStudy (anonymous):

you cant because its the product of two functions i used the chain rule to differentiate e^(-ct) = -c * e^(-ct)